A curve


y  ax  bx  c

2


where a, b and c are constants, passes


through the points (2,11), (-1,-16) and (3,28).

(a) By using the above information, construct a system

containing three linear equations.

(b) Express the above system as a matrix equation


AX  B.

(c) Find the inverse of matrix A by using the adjoint matrix

method. Hence, obtain the values of a, b and c.


1
Expert's answer
2021-09-06T14:29:01-0400
"y=ax^2+bx+c"

(a)


"a(2)^2+b(2)+c=11"

"a(-1)^2+b(-1)+c=-16"




"a(3)^2+b(3)+c=28"

Construct a system containing three linear equations.


"4a+2b+c=11""a-b+c=-16""9a+3b+c=28"

(b)


"AX=B"


"A=\begin{pmatrix}
 4 & 2 & 1 \\
 1 & -1 & 1 \\
 9 & 3 & 1 \\
\end{pmatrix}, X=\begin{pmatrix}
 a \\
 b \\
 c
\end{pmatrix}, B=\begin{pmatrix}
11\\
 -16 \\
 28
\end{pmatrix}"

"\begin{pmatrix}
 4 & 2 & 1 \\
 1 & -1 & 1 \\
 9 & 3 & 1 \\
\end{pmatrix}\begin{pmatrix}
 a \\
 b \\
 c
\end{pmatrix}=\begin{pmatrix}
11\\
 -16 \\
 28
\end{pmatrix}"

(c)


"AX=B"

"A^{-1}AX=A^{-1}B"

"X=A^{-1}B"

"\det A=\begin{vmatrix}
 4 & 2 & 1 \\
 1 & -1 & 1 \\
 9 & 3 & 1 \\
\end{vmatrix}=4\begin{vmatrix}
 -1 & 1 \\
 3 & 1
\end{vmatrix}-2\begin{vmatrix}
 1 & 1 \\
 9 & 1
\end{vmatrix}+1\begin{vmatrix}
 1 & -1 \\
 9 & 3
\end{vmatrix}"

"=4(-1-3)-2(1-9)+(3+9)"

"=-16+16+12=12\not=0=>A^{-1}\ exists"


Find the cofactor matrix:


"C_{11}=(-1)^{1+1}\begin{vmatrix}
 -1 & 1 \\
 3 & 1
\end{vmatrix}=-4"

"C_{12}=(-1)^{1+2}\begin{vmatrix}
 1 & 1 \\
 9 & 1
\end{vmatrix}=8"

"C_{13}=(-1)^{1+3}\begin{vmatrix}
 1 & -1 \\
 9 & 3
\end{vmatrix}=12"

"C_{21}=(-1)^{2+1}\begin{vmatrix}
 2 & 1 \\
 3 & 1
\end{vmatrix}=1"

"C_{22}(-1)^{2+2}\begin{vmatrix}
 4 & 1 \\
 9 & 1
\end{vmatrix}=-5"

"C_{23}=(-1)^{2+3}\begin{vmatrix}
 4 & 2 \\
 9 & 3
\end{vmatrix}=6"

"C_{31}=(-1)^{3+1}\begin{vmatrix}
 2 & 1 \\
 -1 & 1
\end{vmatrix}=3"

"C_{32}=(-1)^{3+2}\begin{vmatrix}
 4 & 1 \\
 1 & 1
\end{vmatrix}=-3"

"C_{33}=(-1)^{3+3}\begin{vmatrix}
 4 & 2 \\
 1 & -1
\end{vmatrix}=-6"

The cofactor matrix is


"C=\begin{pmatrix}
 -4 &8 & 12 \\
 1 & -5 & 6 \\
 3 & -3 & -6 \\
\end{pmatrix}"

The transpose of the cofactor matrix is


"C^T=\begin{pmatrix}
 -4 & 1 & 3\\
 8 & -5 & -3 \\
 12 & 6 & -6 \\
\end{pmatrix}"

"A^{-1}=\dfrac{1}{12}\begin{pmatrix}
 -4 & 1 & 3\\
 8 & -5 & -3 \\
 12 & 6 & -6 \\
\end{pmatrix}"

"A^{-1}=\begin{pmatrix}
 -1/3 & 1/12 & 1/4\\
 2/3 & -5/12 & -1/4 \\
 1 & 1 /2& -1/2 \\
\end{pmatrix}"

"\begin{pmatrix}
 a \\
 b \\
 c
\end{pmatrix}=\begin{pmatrix}
 -1/3 & 1/12 & 1/4\\
 2/3 & -5/12 & -1/4 \\
 1 & 1 /2& -1/2 \\
\end{pmatrix}\begin{pmatrix}
11\\
 -16 \\
 28
\end{pmatrix}"

"=\begin{pmatrix}
-11/3-4/3+7\\
22/3+20/3-7 \\
 11-8-14
\end{pmatrix}=\begin{pmatrix}
2\\
7 \\
 -11
\end{pmatrix}"

"a=2, b=7, c=-11"

"y=2x^2+7x-11"